[figure: 9.7 cm]

THE HALL EFFECT IN Co-RICH THIN FILMS

R. Dlugoš^{1)}, P. Matta^{1)}, A. Zorkovská^{1)}, J. Balcerski^{2)}, P. Sovák^{1)} 1)Faculty of Science, P.J.Šafárik University, Park Angelinum 9, 041 54 Košice, Slovakia 2)Department of Solid State Physics, University of Łódź, Pomorska 141, Łódź, Poland

The aim of this work was to study the influence of structural changes on Hall effect in Co-M-Si-B (M = Fe, Mn, Ni) thin films after annealing. Thin films were prepared by ion-beam sputtering. Crystallographic structure was observed by TEM in as-deposited state and also after annealing in a vacuum furnace at temperature range 300-500^{o}C. Hall effect was measured using DC method. Structural changes in the samples after annealing caused expected decreasing of spontaneous Hall coefficient of all samples with M=Mn,Ni in whole range of annealing temperatures. On the other hand, two ferromagnetic phases could be detected from Hall curves ρ_{H}(B) measured on samples with M=Fe annealed at temperatures 400-450^{o}C (fig.). After annealing at temperature of 500^{o}C unconventional shape of ρ_{H}(B) curve was observed for all the samples which is probably caused by presence of columnar Co crystals with uniaxial anisotropy perpendicular to the surface of sample.

[figure: 13.4 cm]

Subject category :

2. Magnetic Films, Surfaces and Multilayers

Presentation mode : poster

Corresponding author : Rastislav Dlugoš

Address for correspondence : Department of Experimental Physics Faculty of Sciences P.J.Šafárik University Park Angelinum 9 041 54 Košice Slovak Republic

Email address : dlugos@kosice.upjs.sk